Solve these quadratic equations by factoring.
4x2-25=0
-9x2+12x-4=0
6=x2-x

Answer by checkley77(12844)

4x^2-25=0
(2X+5)(2X-5) ANS.
-----------------------
-9x^2+12x-4=0
9X^2-12X+4=0
(3X-2)^2=0 ANS.
----------------------
6=x^2-x
X^2-X-6=0
(X-3)(X+2) ANS.